OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

OpenStudy (anonymous):

Here's the formula to find the slope:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

You should be able to solve it from there.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

You really have no idea how to plug your values in? Really?

OpenStudy (anonymous):

Didn't they teach you this? I'd imagine they taught you if they're giving you this question.

OpenStudy (anonymous):

slope is "m" in the following formula:\[m = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]where you have the given points:\[(x _{1}, y _{1})\]and \[(x _{2}, y _{2})\]which are: (15, 45) and (150, 60) Substitute and simplify. You should be able to match up the variables and points from this.

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

Sigh

OpenStudy (anonymous):

What are you confused about specifically?

OpenStudy (anonymous):

everything lol

OpenStudy (anonymous):

I'm not about to give you the answer, I hope you know that. Here though, I'll plug your values in:\[\frac{ 60 - 45 }{ 150 - 15 }\] Please tell me you know how to do that... and that you see what I did to plug the values in....

OpenStudy (anonymous):

Try to identify your points in regards to x1, y1, x2, y2. In our posts, we have given a one-to-one correspondence. You really wouldn't even need to know algebra to finish it off now. Just substitute. All you have to do is read our helps. It's all there. I can't imagine giving any further help without just giving away an answer, which we are not allowed to do.

OpenStudy (anonymous):

15/135??

OpenStudy (anonymous):

What's with the question mark?

OpenStudy (anonymous):

I'm asking if its right

OpenStudy (anonymous):

Correct. Do you see how I plugged the values in though? And yes, don't question yourself when giving an answer.

OpenStudy (anonymous):

Also, you can simplify 15/135 I think.

OpenStudy (anonymous):

yes thank you very much

OpenStudy (anonymous):

uw.